OpenStudy (anonymous):

PLEASE HELP ME (2x+7/x^2-25) - (1/x-5) = (x^2-4/x^2-25)

OpenStudy (anonymous):

LCD- least common denominator

OpenStudy (compassionate):

Hello, my name is Stephen and I'll be your tutor. \[\frac{ 2x + 7 }{ x^2 - 25 } - \frac{ 1 }{ x - 5 } = \frac{ x^2 - 4 }{ x^2 - 25}\] The first step is finding a common denominator between all of these. The lowest one will do. That way, you can cancel out the denominator and only be left with the non-fraction part to make it simple. Your far left and right terms have a common denominator, what can you multiple x - 5 by to get \[x^2 - 5\](remember, what you do to the denominator you must also do to the numerator.

OpenStudy (anonymous):

Thank you so much !

OpenStudy (compassionate):

You're welcome. Come back to OpenStudy again! We're anxious to help you.